OpenStudy (anonymous):

@ AnthonyAnthony What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? 4x - y = 13 4x + y = 13 4x + y = -13

OpenStudy (anonymous):

So , what's the question ?

OpenStudy (anonymous):

What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0? 4x - y = 13 4x + y = 13 4x + y = -13

OpenStudy (anonymous):

Now the web is working of me .

OpenStudy (anonymous):

The equation of a line is typically written as y=mx+b where m is the slope and b is the y-intercept. If you know two points that a line passes through, this page will show you how to find the equation of the line.

OpenStudy (anonymous):

*for

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

ok so you know that the line passes through (4,3) and we know that the equation formula is y=mx+b so what you need to do is plug into the formula

OpenStudy (anonymous):

i think it's 4x + y = 13

OpenStudy (anonymous):

Thanks. That the answer that I was thinking too, just wanted to be sure it was right. Can you help me more?

OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 5 y + 3 = 2x 2x + 2x - 3 = 5 2x + 3 - 2x = 5 2x + 2x + 3 = 5